OpenStudy (anonymous):

Thanks for answering my question about Joanie owing her brother $5. :D XD Joanie owes her brother $5 and she spent $20 for 5 Honey buns. How

OpenStudy (bahrom7893):

20+5=$25

OpenStudy (bahrom7893):

Suppose x is the amount she made on her job. After paying 5 bucks from it, she was left with $20: x-5=20 x=20+5=25

OpenStudy (anonymous):

Thanks